OpenStudy (hockeychick23):

A box plot is drawn using a 12-inch ruler as the scale. The box extends from 3.5 inches to 6.0 inches. Which of the following represents the upper boundary for outliers (that is, the smallest value beyond which you will find outliers)? A. 7.25 inches B. 6 inches C. 9.75 inches D. 10.5 inches E. 3.75 inches

OpenStudy (rational):

can you find the IQR from the given info

OpenStudy (hockeychick23):

ok, would it be 4.75?

OpenStudy (rational):

Nope, IQR = width of the box = 6-3.5 = ?

OpenStudy (hockeychick23):

oh sorry ok its 2.5

OpenStudy (rational):

Yes! work the upper boundary : \[\text{Q3} + 1.5\times \text{IQR}\]

OpenStudy (hockeychick23):

ok so 6 + (1.5*2.5)= 3.75+6= 9.75

OpenStudy (rational):

Yep!

OpenStudy (hockeychick23):

thanks!

OpenStudy (rational):

np